9c^2-121 factor completely. I have no idea where to start.
Found 2 solutions by scott8148, RAY100:
Answer by scott8148(6628)

this is the difference of two squares ___ two perfect squares with a minus in between

the square root of 9c^2 is 3c

the square root of 121 is 11

the factors are (3c+11)(3c-11)

the difference in the signs results in the middle term of the FOILed product being zero (try it!)

Answer by RAY100(1637)

factoring is recognition of patterns
a^2-B^2 = (a+b) (a-b)
9C^2 -121= (3c)^2 -(11)^2= (3c+11)(3c-11)
checking by foil
(3c+11)(3c-11)
9c^2 -33c+33c-121= 9c^2-121